The equation of the tangent plane to h(x,y)=yln(x) at the point (1,5) is
Select one:
a. z=1/5x+1/5y

b. z=x+y


c. z=1/4x+1/4y


d. z=4(x−1)


e. z=5(x+1)


f. z=5(x−1)
1
Expert's answer
2020-05-29T16:09:28-0400

z = h(x0, y0) + h'x(x0, y0)(x-x0) + h'y(x0, y0)(y-y0) = 0 + 5(x-1) + 0(y-5) = 5(x-1)

The correct answer is f.